CO LINESHAPES: A COMPARISON BETWEEN AB INITIO CALCULATIONS AND HIGH-RESOLUTION MEASUREMENTS

\underline{RICK WEHR}, A. VITCU, A.D. MAY, and J.R. DRUMMOND, Dept. of Physics, University of Toronto, Toronto, Ontario, Canada, M5S 1A7; R. CIURYŁO, Instytut Fizyki, Universytet Mikołaja Kopernika, Grudzia̧dzka 5/7, 87-100 Toruń, Poland; F. THIBAULT, UMR 6627 du CNRS, Universite de Rennes, Campus de Beaulieu, 34042 Rennes Cedex, France; L. ZHAN and W.-K. LIU, Dept. of Physics, University of Waterloo, Waterloo, Ontario, Canada, N2L 3G1.

The role of inelastic collisions in the formation of infrared spectral line shapes is explored through a comparison of spectroscopic measurements and ab initio calculations. The shapes of a set of CO-Ar spectral lines in the fundamental band were recorded between −50^{∘}C and +25^{∘}C, and between 0.05 atm and 1 atm, by a difference-frequency laser spectrometer with a resolution of less than 2 MHz and a signal-to-noise ratio greater than 4000:1. The ab initio line shape calculations used broadening coefficients obtained from fully quantal close-coupled calculations, which were in turn based on an assumed molecular potential energy surface. It is found that a failure to distinguish elastic and inelastic collisions in the CO-Ar system leads to incorrect line shape calculations at low pressures.